A fruit juice franchise company has a policy of opening new fruit juice stand only on those areas that have a mean household income of at least ₱30,500 a month. The company is currently considering an area in which to open a new fruit juice stand. The company’s research department took a sample of 25 households from this area and found that the mean monthly income of these households is ₱32,600. Using 5% significance level, would you conclude that the company should open a fruit juice stand in the area? Also, find the 95% confidence interval of the true mean.

1
Expert's answer
2022-05-09T16:14:32-0400

Let standard deviation is "s=4500."

We need to construct the "95\%" confidence interval for the population mean "\mu."

The critical value for "\alpha = 0.05" and "df = n-1 = 24" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.063899."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(32600-2.063899\times\dfrac{4500}{\sqrt{25}},"

"32600+2.063899\times\dfrac{4500}{\sqrt{25}})"

"=(30742.4909,34457.5091)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "30742.4909 < \mu < 34457.5091," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(30742.4909, 34457.5091)."


The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge30500"

"H_a:\mu<30500"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=24" degrees of freedom, and the critical value for a left-tailed test is "t_c = -1.710882."

The rejection region for this left-tailed test is "R = \{t: t < -1.710882\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{32600-30500}{4500/\sqrt{25}}\approx2.3333"

Since it is observed that "t = 2.3333 \ge -1.711=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed test, "df=24" degrees of freedom, "t=2.3333" is "p = 0.98583," and since "p = 0.98583 \ge 0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 30500, at the "\alpha = 0.05" significance level.